OpenStudy (anonymous):

-3n-1=2-3n-8+5 Help? Is N supposed to be 1.. the -3's cancel each other out.. Thanks.

OpenStudy (mathmale):

What's your goal here? Please share the instructions that came with the problem. Thx. Note: You don't have " -3 's " to cancel out. Rather, you have ' -3n ' s ' to cancel out. Hint: What is 2-8+5?

OpenStudy (anonymous):

It was just to sole each expression.

OpenStudy (anonymous):

solve*

OpenStudy (mathmale):

"To solve" means to find the value of the variable that makes the equation true.

OpenStudy (mathmale):

You have: -3n-1=2-3n-8+5 -3n on the left side equals -3n on the right side, so you may cancel out those terms. type out what you have left.

OpenStudy (anonymous):

-1=2-8+5 right?

OpenStudy (mathmale):

Yes. Now combine the 2 and the -8. Type out the entire equation, showing your latest version.

OpenStudy (anonymous):

-1=-1

OpenStudy (mathmale):

Right -1 is always equal to -1. This tells you that your equation is an identity: It's always true, regardless of the value you give to the variable n.

OpenStudy (mathmale):

Reread the original question and determine what you were supposed to do. How would you answer this question?

OpenStudy (anonymous):

So n doesn't matter then?

OpenStudy (mathmale):

You could put it that way. I would much rather that you say "this equation is an identity b ecause it is true for any value of n."